# Automated Privacy Information Annotation in Large Language Model Interactions

### v1

@@ 5.1 Setups
Dataset Splitting. We split the English data at an 8:2 ratio to obtain the training and test set and follow the dataset splits from the original data sources for Chinese data. The detailed training and test set statistics are shown in Appendix E. During the training process, we balance the number of samples with and without privacy leakage in the training data, ensuring a ratio of approximately 1:1.
@@
Models. We take the instruction-tuned version of different series of language models with various scales, including Llama-3.2-1B and 3B llama3 , and Qwen-2.5-1.5B, 3B, 7B, and 72B qwen2 . All the pretrained checkpoints are loaded from huggingface [^2] https://huggingface.co.
@@
Implementation Details. We implement the Privacy Detection Model and all the baselines in PyTorch. The workstation has 8 NVIDIA V100 32G GPUs. To improve training efficiency, we adopt LoRA lora for tuning our model. More training settings are listed in Appendix H.
@@
Figure 6: Performance of different privacy detection baselines with information-level metrics (%).
5.2 Local Privacy Detection Baselines
@@
Tuning-free Baselines Tuning-free methods leverage the inherent capabilities of pre-trained models for privacy detection. (1) Zero-shot Generation (ZG) gpt3 . We directly prompt local LLMs to determine whether the query leaks privacy or to extract privacy phrase and corresponding information. In the prompt, we explicitly specify that privacy information in user dialogue includes various aspects such as opinions, preferences, intentions, etc. We also require that the recognized privacy be relevant to the user and ask the model to think step by step. Finally, we guide the model to output structured results in the specified format. (2) In-Context Learning (ICL) icl . Based on the prompts of ZG, we further insert five examples from training data based on the template for inference with local privacy detection LLM. After the examples, the current user query is added for privacy detection. All corresponding prompts are provided in Appendix G.
@@
Tuning-based Baseline. We use Supervised Fine-Tuning (SFT) with instructions to enable the model to judge privacy leakage or to generate structured phrases and corresponding information based on the given user’s conversational content. Additionally, we find that overly long instruction templates will reduce the model’s sensitivity to user queries, resulting in a performance drop. The instruction template used during training is provided in Appendix G.
5.3 Main Results
@@
Privacy Leakage Classification. We first evaluate the performance of different baselines on the binary classification task and the results are presented in Table 3. We can find that model performance generally improves with an increase in parameter size. Notably, the largest model in our evaluation, Qwen2.5-72B, excels with ZG, achieving 14.08% and 18.24% improvement for query-level accuracy compared to Qwen2.5-1.5B in English and Chinese dataset, respectively. By comparing SFT with ZG and ICL across different models, we can find that the fine-tuned Qwen2.5-7B achieves the best performance on both English and Chinese datasets. Specifically, SFT with Qwen2.5-7B outperforms ZG and ICL with the same model by over 14.59% and 15.35%, respectively.
@@
Privacy Phrase Extraction. We then investigate the quality of extracted privacy phrases by different baselines. From Table 3, we can also observe that ICL outperforms ZG in most cases. For example, with the Qwen2.5-7B, ICL achieves 3.66% and 3.82% improvement for phrase-level F1 over ZG with the same model in English and Chinese dataset, respectively. These results suggest that inserting examples from the training set enhances privacy detection. Additionally, by fine-tuning on our dataset, the 1B model outperforms the direct inference results of the 72B model by 7.03% and 5.57% in English and Chinese dataset respectively, which is friendly for local devices. Furthermore, the fine-tuned Qwen2.5-7B achieves an improvement of 13.68% and 13.25% in phrase-level F1 compared to Qwen2.5-72B with ICL in English and Chinese dataset, respectively. These significant gains underscore the importance of tuning on our custom dataset.
@@
Privacy Information Summarization. Figure 6 presents the performance of different local privacy detection baselines for information-level F1. We can find that by supervised fine-tuning with the training set, Qwen2.5-7B achieves the best performance on different datasets. Specifically, the fine-tuned Qwen2.5-7B outperforms ZG with Qwen2.5-7B by 21.06% and 28.03%, and outperforms ICL with Qwen2.5-7B by 16.80% and 20.33% in English and Chinese dataset, respectively. We can also observe that fine-tuned 1B model outperforms both ZG and ICL with 72B model by over 10.83% and 7.18% in English and Chinese dataset, respectively. However, there still a gap between the performance of more light-weight LLMs and the ground truth in fine-grained privacy extraction. This emphasizes the need for exploring more effective methods grounded in our dataset. More detailed discussion is provided in Appendix I. Furthermore, to explore the real-time inference efficiency of the local fine-tuned models, we measured the average privacy detection time per query in Appendix J.
5.4 Case Study
@@
Table 4: Case study with Qwen2.5-7B-Instruct.
Query :( I am worried I will screw up the computer because i am bad on coding.
ZG Phrase 1: I am bad on coding
Info 1: The user is worried about their coding abilities.
ICL Phrase 1: bad on coding
Info 1: The user expresses concern about their coding abilities and fears making mistakes that could damage a computer.
SFT Phrase 1: I will screw up the computer
Info 1: The user worries about making mistakes with the computer.
Phrase 2: I am bad on coding
Info 2: The user perceives themselves as not skilled in coding.
@@
To better assess the performance of various local privacy detection methods, we present a case study of phrase and information extraction in Table 4. We can observe that both ZG and ICL fail to capture the user’s specific concerns, while ICL generates more detailed privacy information compared to ZG. In contrast, SFT successfully identifies all relevant privacy phrases from the user input and generates accurate privacy information. We also present a bad case for further analysis of these baselines’ performance in Appendix K.
6 Conclusion, Impact and Future Work
@@
In this work, we have studied privacy detection for real-name LLM interaction. We have constructed the first large-scale multilingual dialogue dataset for different privacy detection tasks, achieved through an automated scalable privacy annotation pipeline for privacy phrases extraction and corresponding information annotation. Additionally, we have developed evaluation metrics at query-level, phrase-level, and information-level. Finally, we have evaluated various local privacy detection methods, revealing that privacy detection in LLM interactions remains a challenging task. As for future research opportunity, the dataset provides a valuable foundation to explore more effective privacy detection methods that can be efficiently deployed on different local devices.
Limitations and Discussion. We acknowledge that the built dataset is designed for privacy detection of user-related information in user’s daily interactions with LLM and should not be regarded as a comprehensive dataset, as it may not suitable for all possible privacy leakage scenarios. For example, detecting privacy from user’s query whether it leaks any information from enterprise database. Extending the dataset by applying the built automated annotation pipeline to different private scenarios remains a future prospect. Ethical statement is provided in Appendix M.
@@ Appendix A Prompts for Automated Privacy Annotation Pipeline
Table 10, Table 11 - 12, Table 13 - 16, and Table 17 provides prompts for privacy leakage classification, extensive privacy categories extraction, privacy phrase extraction, and privacy information annotation in designed pipeline, respectively.
Appendix B Introduction of Data Source
ShareGPT sharegpt is composed of dialogues between users and ChatGPT, with 79K raw dialogues and 243K user queries. CrossWOZ crosswoz is a task-oriented dialogue dataset covering five domains, containing 6K dialogues and 47K user queries. DuConv duconv focuses on movie and film star topics, with 22K dialogues and 99K user queries. Both CrossWOZ and DuConv data are collected through dialogues conducted by two workers, based on specific topics or knowledge backgrounds. The LCCC-base lccc is derived from the Weibo corpus. We selected a subset of this dataset as our source data, including 50K dialogues and 76K user queries.
Appendix C Dataset Examples
Table 18 shows raw data examples and examples of the constructed dataset. One data sample is consists of a user-initiated query and a response from the assistant. Based on this, we add privacy detection results to each query, in the form of a list of privacy. Each element in the list consists of a phrase extracted from the query and the corresponding specific privacy information. In terms of dataset processing, we also retain the contextual relationship between the original queries in one dialogue.
Table 5: The comparison with existing dataset.
Dataset Target Labeling Data Size Information Language
I2B2/UTHealth i2b2 Personal Health Manual 1,304 $\times$ EN
CodE ALLTAG codealltag Email Manual 2,309 $\times$ DE
TAB tab Court Judgments Manual 1,268 $\times$ EN
Reddit-SD selfdisclosures Comments Manual 2,415 $\times$ EN
Ours LLM Interactions Automated 249,683 $\surd$ EN, ZH
Appendix D Comparison with Existing Datasets
As shown in Table 5, our dataset is centered on user interactions, where the definition of privacy extends beyond identifiable details and encompasses a broader range of sensitive information. Additionally, our dataset is annotated automatically, resulting in significantly larger scale compared to existing datasets. Furthermore, it offers detailed information for each extracted privacy phrase.
Appendix E Details on Training and Testing Sets
Table 6: Statistics of the dataset with automatically annotated privacy phrases and information.
@@ Appendix E Details on Training and Testing Sets
Table 6 lists detailed statistics of the dialogue dataset with annotated privacy phrases and information. The privacy phrase set and information set is empty for samples without privacy leakage, while the average length of phrase set and information set for samples with privacy leakage is 2.6 in English part and 1.6 in Chinese part.
Appendix F Human Annotation Guidelines
F.1 Annotation Instructions for Privacy Leakage
F.2 Annotation Instructions for Privacy Phrase Extraction
Appendix G Prompts for Local Privacy Detection Methods
Table 19 - Table 21, Table 22 - 24, and Table 25 - 27 shows detailed prompts used for zero-shot generation, in-context learning, and training privacy detection models, respectively. For zero-shot generation and in-context learning, we prompt language models to extract privacy phrases from user dialogues and generate specific information for each identified phrase. For fine-tuning, we employ a concise template to guide the model, allowing it to learn both the content and output format from the training data.
Appendix H Additional Training Details
Appendix I Discussion of Main Results
The main purpose of local privacy detection methods is to perform privacy leakage classification, phrase extraction, or information summarization without the need to upload raw user queries online. Although the fine-tuned Qwen2.5-7B model achieves an accuracy of 87.56% for binary privacy leakage classification task, its performance on more complex tasks can not meet the requirements for practical implementation. For example, the phrase-level F1 score for the best fine-tuned model is 74.27% and the best information-level F1 score is 44.70% on the Chinese dataset. This also highlights a gap between the performance of more light-weight LLMs and the ground truth in fine-grained privacy extraction. To address this, a research line is to explore new methods to enhance the performance of generative locally deployable LLMs on these tasks, including the design of novel model architectures and the incorporation of different training methods, such as reinforcement learning.
Appendix J Inference Time for Local Privacy Detection Models
Table 7: Average privacy detection time per query (s) of fine-tuned local models for English samples.
@@ Appendix I Discussion of Main Results
Table 8: Average privacy detection time per query (s) of fine-tuned local models for Chinese samples.
@@
To explore the real-time inference efficiency of the local privacy detection models which are fine-tuned on our dataset, we measured the average privacy detection time per query for both English and Chinese samples. The model precision is set to FP16 and the results is shown in Table 7 and Table 8. These results reveal the time efficiency of local privacy detection with billion-scale LLMs.
Appendix K Case Study
Table 9: A bad case of Qwen2.5-7B-Instruct with SFT.
Query My 3 year old repeats syllables twice, rather than pronouncing both syllables of the word. For example he says ‘The taitai needs wawa’ instead of ‘the tiger needs water.’ He sounds a bit babyish. Why does he do this? How can I help him? Should I be worried about it?
SFT Phrase 1: repeats syllables twice
Information 1: The user is concerned about their 3-year-old child’s speech development, specifically regarding the pronunciation of syllables.
Phrase 2: Should I be worried about it?
Information 2: The user is concerned about their 3-year-old child’s speech development.
GT Phrase 1: My 3 year old
information 1: The user’s child is 3 years old.
Phrase 2: He sounds a bit babyish.
Information 2: The user’s child is described as having speech that sounds babyish, which suggests a concern about speech development.
Phrase 3: repeats syllables twice
Information 3: The user’s 3-year-old child has a speech pattern of repeating syllables twice rather than pronouncing both syllables of words.
@@
Current local privacy detection methods also have some limitations. Table 9 presents a bad case from the fine-tuned Qwen2.5-7B, where the local model incorrectly extracted a general question that did not actually expose the user’s privacy. Moreover, the model failed to identify certain implicit privacy phrases. These shortcomings highlight the challenges of extracting phrases and generating accurate information. Light-weight LLMs, with limited text comprehension abilities, require more sophisticated methods to be developed in future work for effective privacy detection on the user side.
Appendix L API Cost
@@
We use gpt-4o-2024-08-06 as the base LLM for our automated privacy annotation pipeline, which extracts privacy phrases and annotates information. For the English dataset, the total token count is 296M tokens for input and 152M for output. For the Chinese dataset, the total token count is 219M tokens for input and 99M for output. The total annotation cost amounts to approximately \$3,800 and the average cost is \$0.015 per sample.
Appendix M Ethical Statement
@@
We investigates privacy detection in user’s real-name interactions with LLM. The design, implementation, and use of the automated privacy annotation pipeline, along with the local privacy detection methods, are guided by ethical principles for responsible AI deployment. (1) Dataset Construction: All dialogue data used in this study are publicly available and properly cited, with no ethical concerns regarding their use for research purposes. (2) Fair Usage: The primary goal of the privacy annotation pipeline is to extract privacy phrases and annotate privacy information from public dialogue datasets, rather than conducting malicious analysis of user privacy. We believe that the responsible application of our design aligns with ethical guidelines.
Table 10: Prompt template for privacy leakage classification for English and Chinese dataset, where <|INPUT|> is the input of user query.
@@
Prompts for privacy leakage classification (English).
[…]
A user poses a query
"<|INPUT|>"
@@
Give your think steps. Your reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
@@
“‘json
    "reason": reason for your judgment,
    "judgment": true or false
Here are some examples: User’s query: We want to play football today. JSON Output: “‘json
    "reason": "This query exposes the activity of the user and people related to the user intend to do.",
    "judgment": true
…… (other examples)
Prompts for privacy leakage classification (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn
{CJK}UTF8gbsn 在用户提出请求之前,你只知道用户的真实的唯一标识符,并且不知道用户的任何其他的隐私敏感信息。
{CJK}UTF8gbsn 在收到带有用户真实的唯一标识符的请求后,请你进行逐步的思考(Think Step By Step)来判断该请求是否向你透露了任何新的关于用户或与用户相关的人或事的隐私敏感信息
{CJK}UTF8gbsn
{CJK}UTF8gbsn 隐私敏感信息应与用户或与用户相关的人或事相关,该相关性体现在用户请求的上下文中。
{CJK}UTF8gbsn 隐私敏感信息包括任何观点、担忧、偏好、活动、意图或任何其他根据查询上下文可被视为私密或敏感的信息。
{CJK}UTF8gbsn
{CJK}UTF8gbsn 给出你的思考与分析步骤。你的推理必须依赖于查询中包含的直接证据,不能基于假设
{CJK}UTF8gbsn
{CJK}UTF8gbsn 最后,输出如下格式的json:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason": 你判断的理由,
{CJK}UTF8gbsn     "judgment": true 或 false
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:我们今天想踢足球。
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason": "该请求透露了用户及与用户相关的人打算进行的活动。",
{CJK}UTF8gbsn     "judgment": true
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn …… (other examples)
Table 11: Prompt template for privacy categories extraction in step 2 for English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for privacy categories extraction (English).
[…]
Given a user’s query
"<|INPUT|>"
Identify which phrases in the query leak the user’s privacy information and the correponding categories of privacy information.
@@
“‘json
    privacy information category 1: [phrase 1, phrase 3, …],
    privacy information category 2: [phrase 2, phrase 4, …]
Here is an example
"Query": "I’m a teacher at China, majoring in computer science. I will go back home to London."
@@
“‘json
    "Job": ["teacher"],
    "Subject": ["computer science"],
    "Location": ["China", "London"]
Prompts for privacy categories extraction (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn
{CJK}UTF8gbsn 提取请求中泄露用户隐私信息的短语,并将其分类为相应的隐私信息类别
{CJK}UTF8gbsn 让我们一步一步地思考,最后输出一个json,格式如下:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "隐私信息类别1": ["短语1", "短语3", …],
{CJK}UTF8gbsn     "隐私信息类别2": ["短语2", "短语4", …]
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一个示例:
{CJK}UTF8gbsn 用户的请求:"我是中国的一名教师,主修计算机科学。我将回家到伦敦。"
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "职业": ["教师"],
{CJK}UTF8gbsn     "专业": ["计算机科学"],
{CJK}UTF8gbsn     "地点": ["中国", "伦敦"]
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
Table 12: Prompt template for categories deduplication in step 2 for English and Chinese dataset, where <|INPUT|> is the input of privacy categories and corresponding phrase examples, and <|BLOCK-SIZE|> is the number of input categories.
Prompts for categories deduplication (English).
[…]
Given the following <|BLOCK-SIZE|> privacy categories and the list of example phrases belong to each privacy category:
<|INPUT|>

### v2

@@ 5.1. Setups
Dataset Splitting. We split the English data at an 8:2 ratio to obtain the training and test set and follow the dataset splits from the original data sources for Chinese data. The detailed training and test set statistics are shown in Appendix B.3. During the training process, we balance the number of samples with and without privacy leakage in the training data, ensuring a ratio of approximately 1:1.
@@
Models. We take the instruction-tuned version of different series of language models with various scales, including Llama-3.2-1B and 3B (Dubey et al., 2024), and Qwen-2.5-1.5B, 3B, 7B, and 72B (Yang et al., 2024). All the pretrained checkpoints are loaded from huggingface [^2] https://huggingface.co.
@@
Implementation Details. We implement all the baselines in PyTorch. The workstation has 8 NVIDIA V100 32G GPUs. To improve training efficiency, we adopt LoRA (Hu et al., 2022) for tuning our model. More training settings are listed in Appendix D.1.
@@
Figure 6. Performance of different privacy information summarization baselines from information-level F1 (%).
5.2. Local Privacy Detection Baselines
@@
Tuning-free Baselines Tuning-free methods leverage the inherent capabilities of pre-trained models for privacy detection. (1) Zero-shot Generation (ZG) (Brown et al., 2020). We directly prompt local LLMs to determine whether the query leaks privacy or to extract privacy phrases and corresponding information. In the prompt, we explicitly specify that privacy information in user dialogue includes various aspects such as opinions, preferences, intentions, etc. We also require that the recognized privacy be relevant to the user and ask the model to think step by step. Finally, we guide the model to output structured results in the specified format. (2) In-Context Learning (ICL) (Dong et al., 2023). Based on the prompts of ZG, we further insert five examples randomly sampled from training data based on the template for inference with local privacy detection LLM. After the examples, the current user query is added for privacy detection. All corresponding prompts are provided in Appendix D.2.
@@
Tuning-based Baseline. We use Supervised Fine-Tuning (SFT) with instructions to enable the model to judge privacy leakage or to generate structured phrases and corresponding information based on the given conversations. Additionally, we find that overly long instruction templates will reduce the model’s sensitivity to user queries, resulting in a performance drop. The instruction template used during training is also provided in Appendix D.2.
5.3. Main Results
@@
Privacy Leakage Classification. We first evaluate the performance of different baselines on the binary classification task, and the results are presented in Table 4. We can find that model performance generally improves with an increase in parameter size. Notably, the largest model in our evaluation, Qwen2.5-72B, excels with ZG, achieving 14.08% and 18.24% improvement for query-level accuracy compared to Qwen2.5-1.5B in English and Chinese datasets, respectively. By comparing SFT with ZG and ICL across different models, we can find that the fine-tuned Qwen2.5-7B achieves the best performance on both English and Chinese datasets. Specifically, SFT with Qwen2.5-7B outperforms ZG and ICL with the same model by over 14.59% and 15.35%, respectively.
@@
Privacy Phrase Extraction. We then investigate the quality of extracted privacy phrases by different baselines. From Table 4, we can also observe that ICL outperforms ZG in most cases. For example, with the Qwen2.5-7B, ICL achieves 3.66% and 3.82% improvement for phrase-level F1 over ZG with the same model in English and Chinese datasets, respectively. These results suggest that inserting examples from the training set enhances privacy detection. Additionally, by fine-tuning on our dataset, the 1B model outperforms the direct inference results of the 72B model by 7.03% and 5.57% in English and Chinese datasets, respectively, which is friendly for local devices. Furthermore, the fine-tuned Qwen2.5-7B achieves an improvement of 13.68% and 13.25% for phrase-level F1 compared to Qwen2.5-72B with ICL in English and Chinese datasets, respectively. These significant gains underscore the importance of tuning with our custom dataset.
@@
Privacy Information Summarization. Figure 6 presents the performance of different local privacy detection baselines from information-level F1. We can find that by supervised fine-tuning with the training set, Qwen2.5-7B achieves the best performance on different datasets. Specifically, the fine-tuned Qwen2.5-7B outperforms ZG with Qwen2.5-7B by 21.06% and 28.03%, and outperforms ICL with Qwen2.5-7B by 16.80% and 20.33% in English and Chinese datasets, respectively. We can also observe that the fine-tuned 1B model outperforms both ZG and ICL with the 72B model by over 10.83% and 7.18% in English and Chinese datasets, respectively.
Discussion. Although the fine-tuned Qwen2.5-7B model achieves a binary privacy leakage classification accuracy of 87.56%, its performance on more complex tasks remains insufficient for practical deployment. For instance, the best fine-tuned model attains a phrase-level F1 score of 74.27% and an information-level F1 score of only 44.70% on the Chinese dataset. These results emphasize the need for exploring more effective methods grounded in our dataset. Furthermore, to explore the real-time inference efficiency of the local fine-tuned models, we also measured the average privacy detection time per query in Appendix D.3.
5.4. Comparison with Traditional PII Detection
Table 5. Performance of traditional PII detection methods from phrase-level metrics (%)
Method
English Chinese
$\text{R}_{\mathcal{P}}$
$\text{P}_{\mathcal{P}}$
$\text{F1}_{\mathcal{P}}$
$\text{R}_{\mathcal{P}}$
$\text{P}_{\mathcal{P}}$
$\text{F1}_{\mathcal{P}}$
Presidio (Microsoft, 2020) 13.51 25.15 17.58 19.38 22.64 20.89
Presidio [transformer] 14.73 27.13 19.09 18.52 22.68 20.39
Qwen2.5-7B with SFT 70.75 60.31 65.11 80.15 69.20 74.27
We also evaluate open-source traditional PII detection methods on our datasets, including Presidio with its default spaCy-based model and an alternative configuration using a transformer-based named entity recognition model (Microsoft, 2020). The results, presented in Table 5, show that these traditional methods perform substantially worse than the Qwen2.5-7B model fine-tuned on our dataset. This performance gap also indicates that the nature of privacy information in user interactions with LLMs differs markedly from the PII typically found in structured documents or publicly available web content.
5.5. Case Study
@@
Figure 7. Case study with Qwen2.5-7B-Instruct. SFT successfully detects all relevant privacy details.
@@
To better assess the performance of various local privacy detection methods, we present a case study of phrase and information extraction in Figure 7. We can observe that both ZG and ICL fail to capture the user’s specific concerns, while ICL generates more detailed privacy information compared to ZG. In contrast, SFT successfully identifies all relevant privacy phrases from the user input and generates accurate privacy information. We present a failure case for further analysis of these baselines’ performance in Appendix D.4.
6. Conclusion
@@
In this work, we have studied privacy detection for real-name LLM interaction. We have constructed the first large-scale multilingual dialogue dataset for different privacy detection tasks, achieved through an automated scalable privacy annotation pipeline for privacy phrase extraction and corresponding information annotation. Additionally, we have developed evaluation metrics at the query-level, phrase-level, and information-level. Finally, we have evaluated various local privacy detection methods, revealing that privacy detection in LLM interactions remains a challenging task. As for future research opportunities, the dataset provides a valuable foundation to explore more effective privacy detection methods that can be efficiently deployed on different local devices.
@@ Appendix A Prompts for Automated Privacy Annotation Pipeline
Figure 9, Figure 10 - 11, Figure 12 - 15, and Figure 16 provides prompts for privacy leakage classification, extensive privacy categories extraction, privacy phrase extraction, and privacy information annotation in designed pipeline, respectively.
Appendix B Additional Dataset Details
B.1. Introduction of Data Source
ShareGPT (ShareGPT, 2023) is composed of dialogues between users and ChatGPT, with 79K raw dialogues and 243K user queries. CrossWOZ (Zhu et al., 2020) is a task-oriented dialogue dataset covering five domains, containing 6K dialogues and 47K user queries. DuConv (Wu et al., 2019) focuses on movie and film star topics, with 22K dialogues and 99K user queries. Both CrossWOZ and DuConv data are collected through dialogues conducted by two workers, based on specific topics or knowledge backgrounds. The LCCC-base (Wang et al., 2020) is derived from the Weibo corpus. We selected a subset of this dataset as our source data, including 50K dialogues and 76K user queries.
B.2. Dataset Examples
Figure 17 shows raw data examples and examples of the constructed dataset. One data sample consists of a user-initiated query and a response from the assistant. Based on this, we add privacy detection results to each query, in the form of a list of privacy. Each element in the list consists of a phrase extracted from the query and the corresponding specific privacy information. In terms of dataset processing, we also retain the contextual relationship between the original queries in one dialogue.
Table 6. Statistics of the dataset with automatically annotated privacy phrases and information.
@@ B.2. Dataset Examples
B.3. Details on Training and Testing Sets
Table 6 lists detailed statistics of the dialogue dataset with annotated privacy phrases and information. The privacy phrase set and information set are empty for samples without privacy leakage, while the average length of the phrase set and information set for samples with privacy leakage is 2.6 in the English part and 1.6 in the Chinese part.
Table 7. Average privacy detection time per query (s) of fine-tuned local models for English samples.
@@ B.3. Details on Training and Testing Sets
Table 8. Average privacy detection time per query (s) of fine-tuned local models for Chinese samples.
@@
Appendix C Human Annotation and Evaluation Guidelines
C.1. Annotation Instructions for Privacy Leakage
C.2. Annotation Instructions for Privacy Phrase Extraction
C.3. Evaluation Instructions for Annotated Privacy Information
Annotators are asked to evaluate the annotated privacy information corresponding to each privacy phrase. A score of 1 is assigned if the generated privacy information accurately captures the details of the phrase and is semantically coherent; otherwise, it receives a score of 0. The final score is computed as the average of all individual scores.
Figure 8. A failure case of Qwen2.5-7B-Instruct with SFT.
Appendix D Additional Experiment Details and Results
D.1. Additional Training Details
D.2. Prompts for Local Privacy Detection Methods
Figure 18 - Figure 20, Figure 21 - 23, and Figure 24 - 26 shows detailed prompts used for zero-shot generation, in-context learning, and training privacy detection models, respectively. For zero-shot generation and in-context learning, we prompt language models to extract privacy phrases from user dialogues and generate specific information for each identified phrase. For fine-tuning, we employ a concise template to guide the model, allowing it to learn both the content and output format from the training data.
D.3. Inference Time for Local Privacy Detection Models
@@
To explore the real-time inference efficiency of the local privacy detection models, which are fine-tuned on our dataset, we measured the average privacy detection time per query for both English and Chinese samples. The model precision is set to FP16, and the results are shown in Table 7 and Table 8. These results reveal the time efficiency of local privacy detection with billion-scale LLMs.
D.4. Case Study
@@
Figure 8 presents a failure case from the fine-tuned Qwen2.5-7B, where the local model incorrectly extracts a general question that does not actually expose the user’s privacy. Moreover, the model fails to identify certain implicit privacy phrases. These shortcomings highlight the challenges of extracting phrases and generating accurate information. Light-weight LLMs, with limited text comprehension abilities, require more sophisticated methods to be developed in future work for effective privacy detection on the user side.
Appendix E API Cost
@@
We use gpt-4o-2024-08-06 as the base LLM for our automated privacy annotation pipeline, which extracts privacy phrases and annotates information. For the English dataset, the total token count is 296M tokens for input and 152M for output. For the Chinese dataset, the total token count is 219M tokens for input and 99M for output. The total annotation cost amounts to approximately \$3,800, and the average cost is \$0.015 per sample.
Appendix F Ethical Statement
@@
We investigate privacy detection in users’ real-name interactions with LLM. The design, implementation, and use of the automated privacy annotation pipeline, along with the local privacy detection methods, are guided by ethical principles for responsible AI deployment. (1) Dataset Construction: All dialogue data used in this study are publicly available and properly cited, with no ethical concerns regarding their use for research purposes. (2) Fair Usage: The primary goal of the privacy annotation pipeline is to extract privacy phrases and annotate privacy information from public dialogue datasets, rather than conducting malicious analysis of user privacy. We believe that the responsible application of our design aligns with ethical guidelines. Furthermore, to apply our annotation pipeline to a new corpus, such as private enterprise documents, one can begin by annotating a small subset of the preprocessed data with our pipeline. A privacy detection model can then be trained on this subset to generate labels for the remaining data. Note that all sensitive content uploaded to GPT-4o or any cloud-based LLM should be carefully reviewed to ensure data security and compliance.
@@
Prompts for privacy leakage classification (English). A user poses a query
”¡—INPUT—¿”
@@
Give your think steps. Your reasoning must depend on the direct evidence contained in the query and cannot be based on assumption.
@@
‘‘‘json { "reason": reason for your judgment, "judgment": true or false } ’’’ Here are some examples: User’s query: We want to play football today. JSON Output: ‘‘‘json { "reason": "This query exposes the activity of the user and people related to the user intend to do.", "judgment": true } ’’’ …… (other examples)
Prompts for privacy leakage classification (Chinese). 用户向你提出一个请求
”¡—INPUT—¿”
在用户提出请求之前,你只知道用户的真实的唯一标识符,并且不知道用户的任何其他的隐私敏感信息。
在收到带有用户真实的唯一标识符的请求后,请你进行逐步的思考(Think Step By Step)来判断该请求是否向你透露了任何新的关于用户或与用户相关的人或事的隐私敏感信息
隐私敏感信息应与用户或与用户相关的人或事相关,该相关性体现在用户请求的上下文中。
隐私敏感信息包括任何观点、担忧、偏好、活动、意图或任何其他根据查询上下文可被视为私密或敏感的信息。
给出你的思考与分析步骤。你的推理必须依赖于查询中包含的直接证据,不能基于假设
最后,输出如下格式的json:
‘‘‘json { "reason": 你判断的理由, "judgment": true or false } ’’’ 以下是一些示例:
用户的请求:我们今天想踢足球。
JSON输出:
‘‘‘json { "reason": " 该请求透露了用户及与用户相关的人打算进行的活动。 ", "judgment": true } ’’’ …… (other examples)
Figure 9. Prompt template for privacy leakage classification for English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query.
Prompts for privacy categories extraction (English). Given a user’s query
”¡—INPUT—¿”
Identify which phrases in the query leak the user’s privacy information and the corresponding categories of privacy information.
@@
‘‘‘json { privacy information category 1: [phrase 1, phrase 3, ...], privacy information category 2: [phrase 2, phrase 4, ...] } ’’’ Here is an example
”Query”: ”I’m a teacher in China, majoring in computer science. I will go back home to London.”
@@
‘‘‘json { "Job": ["teacher"], "Subject": ["computer science"], "Location": ["China", "London"] } ’’’
Prompts for privacy categories extraction (Chinese). 用户向你提出一个请求
”¡—INPUT—¿”
提取请求中泄露用户隐私信息的短语,并将其分类为相应的隐私信息类别
让我们一步一步地思考,最后输出一个json,格式如下:
‘‘‘json { "隐私信息类别1" : [ "短语1", "短语3" , ...], "隐私信息类别2" : [ "短语2", "短语4" , ...] } ’’’ 以下是一个示例:
用户的请求:”我是中国的一名教师,主修计算机科学。我将回家到伦敦。”
JSON输出:
‘‘‘json { "职业" : [ "教师" ], "专业" : [ "计算机科学" ], "地点" : [ "中国", "伦敦" ] } ’’’
Figure 10. Prompt template for privacy categories extraction in step 2 for English and Chinese datasets, where ¡—INPUT—¿ is the input of the user query.
Prompts for categories deduplication (English). Given the following ¡—BLOCK-SIZE—¿ privacy categories and the list of example phrases belonging to each privacy category:
¡—INPUT—¿